# MasterWeaver: Taming Editability and Identity for Personalized Text-to-Image Generation

### v2

@@ 3.4 Learning Objective
Following [39], we also employ a background disentanglement loss to encourage that the identity feature only controls the generation of the facial region. Specifically, when the identity feature changes, we regularize the background to be not changed,
@@
where $f=M\circ\psi(x)$ is the extracted identity feature and $1-m$ denotes the mask of background region. $wei$ denotes the augmentation operation.
@@ 4.1 Experimental Settings
Evaluation dataset. To evaluate the one-shot personalization, we randomly select 300 identities from the CelebA-HQ dataset [35], and each identity has one image as the reference. Besides, our method can be extended for personalized generation using multiple reference images without additional training. Here, we also evaluate our method with multiple images as the reference. Following [40], we collect a dataset consisting of 25 identities, and each identity contains four images for evaluation. For quantitative analysis, we employ 50 prompts that encompass a range of clothing, styles, attributes, actions, and backgrounds. We randomly generate five images for each identity-prompt pair. More details are provided in Suppl. For visual comparison, we utilize the identity images collected from existing methods for a convenient comparison. Unless specifically noted, all images presented in this paper are produced using a single reference image.
Implementation details. We employ Stable Diffusion V1.5 in our experiments, and our model is trained using a batch size of 16 and a learning rate of 1e-6. To calculate the $\mathcal{L}_{edit}$, we utilize the output features of the cross attention blocks in the first decoder layer. $\lambda_{edit}$ is set as 0.01 and $\lambda_{disen}$ is set as 1. To enable classifier-free guidance, we use a probability of 0.05 to drop text and image individually, and a probability of 0.05 to drop text and image simultaneously. All experiments are conducted on 4$\times$×A800 GPUs with AdamW [47] optimizer. During image generation, we use 50 steps of the LMS sampler, and the scale of classifier-free guidance is 5. More details of our model and implementation are provided in the Suppl.
@@
Figure 7: Ablation Study. The proposed losses and augmented dataset (denotes as aug) significantly improves the editability of the model. Besides, the identity fidelity consistently improves as the number increases.
@@ 4.2 Qualitative Comparison
Fig. 5 illustrates visual comparisons with baseline methods using a single image as a reference. As depicted, our MasterWeaver can generate photo-realistic personalized images in diverse scenarios, including modifications to attributes, clothing, background, and style. Optimization-based methods, i.e., Textual Inversion, Dreambooth, and Custom Diffusion, usually suffer from the overfitting in this limited-data scenario, leading to limited text alignment and compromised identity fidelity (e.g., rows 1$\sim$3). Tuning-free methods, i.e., Fastcomposer and IP Adapter maintain high identity fidelity but fall short in editing style and facial attributes (e.g., face attributes in rows 1 and 3, and the style in the last two rows). Photomaker shows improved text controllability; however, it sometimes fails to generate faithful identity (e.g., rows 1, 2, and 5). In comparison, our MasterWeaver can generate personalized images that are faithful to both reference identity and text prompts. Fig. 1 illustrates more results generated by our method, and our MasterWeaver can generate photo-realistic images with diverse clothing, accessories, facial attributes, and actions. Additionally, it allows for the simultaneous editing of multiple attributes.
MasterWeaver can also be directly extended for personalized image generation with multiple reference images of the same identity. Specifically, during inference, we simply concatenate the identity features of different images along the token dimension. As shown in Fig. 7, using more reference images improves the identity fidelity of the generated personalized images. Furthermore, we compare our method with baseline methods in a setting with multiple reference images (i.e., with four reference images). From Fig. 6, our MasterWeaver still outperforms competitors in identity fidelity and text controllability. More qualitative results can be found in Suppl.
Table 3: User study. The numbers indicate the percentage (%) of volunteers who favor the results of our method over those of the competing methods.
Text Fidelity 68.8 61.5 62.2 70.2 52.1
Identity Fidelity 64.0 70.1 55.3 53.5 66.8
Image Quality 58.2 74.1 68.8 71.8 54.5
Table 4: Ablation Study. The proposed background disentanglement loss $L_{disen}$, editing direction loss $L_{edit}$, and the face-augmented dataset (denote as aug) can improve the editability of our MasterWeaver, while keeping the identity fidelity.
@@ 4.3 Quantitative Comparison
In addition to the qualitative comparisons, we further conduct the quantitative evaluation to validate the performance of our MasterWeaver. Table 1 reports the comparisons conducted with a single reference image. From the table, our method outperforms existing state-of-the-art methods regarding text alignment and face similarity, demonstrating its ability to generate personalized images that maintain consistent identity and are well-aligned with texts. Moreover, our MasterWeaver exhibits a competitive inference speed, requiring only 4s to generate an image on a single V100 GPU. Table 2 reports similar findings in a scenario utilizing four reference images, further demonstrating the effectiveness of our MasterWeaver. While our scores for CLIP-I and DINO are not as high, it is important to note that these metrics are not specifically designed for facial analysis and can be affected by extraneous factors such as pose. Additionally, we have conducted a user study to compare MasterWeaver with other methods based on human perception.
User Study. Given an identity image, a text prompt, and two synthesized images (ours v.s. competitor’s), the participants are asked to select the better one from three views: i) Text Fidelity: Which image is more faithful to the input text prompt? ii) Identity Fidelity: Which image is most similar to the input person’s identity? iii) Image Quality: Which generated image shows better quality? For each evaluated view, we employ 40 participants, and each participant is asked to answer 50 randomly selected questions, i.e., 2000 responses in total. As shown in Table 3, our method receives better preference than others.
@@ 4.4 Ablation Study
We have conducted the ablation studies to evaluate the effect of various components in our method, including the background disentanglement loss $\mathcal{L}_{disen}$, editing direction loss $\mathcal{L}_{edit}$, and the face-augmented dataset.
Effect of disentanglement loss $\mathcal{L}_{disen}$. We first evaluate the impact of $\mathcal{L}_{disen}$. As illustrated in Fig. 7, without $\mathcal{L}_{disen}$, the model fails to generate images with desired clothing. This suggests that $\mathcal{L}_{disen}$ could ease the influence of identity information on the background, thereby enhancing the editability of the generated images. From Table 4, both text alignment and face similarity metrics decrease without $\mathcal{L}_{disen}$, demonstrating its effectiveness.
Effect of editing direction loss $\mathcal{L}_{edit}$. We then assess the effect of the $\mathcal{L}_{edit}$. As shown in Fig. 7, without the $\mathcal{L}_{edit}$, our MasteWeaver fails to generate the identities with proper facial attributes (i.e., old woman). From Table 4, the CLIP-T score drops significantly without $\mathcal{L}_{edit}$. Although the preservation of identity drops slightly, it is also acceptable.
Effect of the face-augmented dataset. We further evaluate the effect of our constructed face-augmented dataset. Fig. 7 demonstrates that the model trained with the face-augmented dataset exhibits improved editability (e.g., with the augmented dataset, our MasterWeaver successfully generates the personalized image with a smiling expression). The results in Table 4 further confirm its effectiveness, as the text alignment score increases when using the augmented dataset. More ablations are provided in Suppl.
4.5 Limitations
Our method is primarily trained to generate images with a single face, and it may struggle with generating images consisting of multiple personalized identities. Additionally, due to the coarse granularity of text representation, our method also faces challenges in achieving precise control over attributes. A more detailed analysis is available in the Suppl.
@@ 5 Conclusion
In this work, we propose MasterWeaver, a novel tuning-free method capable of generating personalized images with high efficiency, faithful identity, and flexible editability. The proposed editing direction loss and face-augmented dataset significantly improve the model’s editability while maintaining identity fidelity. Extensive experiments have demonstrated that our MasterWeaver outperforms the state-of-the-art methods and can generate photo-realistic images that are faithful to both identity and texts. This capacity makes our method versatile for various applications, including personalized digital content creation and artistic endeavors. Moreover, the proposed editing direction loss has the potential to be applied to other domains (e.g., animals and objects), enhancing its applicability.
@@
Fig. S8 illustrates the detailed framework of our MasterWeaver. As shown in the figure, our ID Mapper consists of a stack of cross attention, feed forward layer and self attention layer. To effectively incorporate the identity feature $f$ with text feature to steer the personalized generation, we adopt the dual cross attention mechanism [87, 82]. As illustrated in Fig. S8, for each cross attention block of stable diffusion, we further introduce two learnable projection layers $W^{f}_{K}$ and $W^{f}_{V}$. The identity feature is then integrated by: $\text{Attention}(Q,K^{f},V^{f})$, where $K^{f}=W^{f}_{K}\cdot f$ and $V^{f}=W^{f}_{V}\cdot f$ represent the projected key and value matrices of the identity information, respectively. To fuse the identity information with the text information, we sum them by:
@@
where $\lambda$ is a trade-off hyperparameter and set as $1$ during training. During training, we optimize the parameters of ID Mapper and the additional introduced projection layers simultaneously, and keep the parameters of stable diffusion fixed.
@@
To calculate the editing direction loss $\mathcal{L}_{edit}$, we utilize the output features from the cross attention blocks in the first decoder layer. Specifically, given the source prompt $y$, target prompt $y^{\prime}$ and reference image $x$, denote the editing directions of the $l$-th cross attention block for stable diffusion and MasterWeaver are $\Delta^{l}_{\epsilon_{\theta}}(y,y^{\prime})\in\mathbb{R}^{H\times W\times D}$ and $\Delta^{l}_{\epsilon_{\theta^{\prime}}}(y,y^{\prime},x)\in\mathbb{R}^{H\times W\times D}$. $H$, $W$, and $D$ represent the height, width, and dimension of the output feature, respectively. Then, the editing direction loss is calculated as follows:
@@
$\mathcal{L}^{l}_{edit}=\frac{1}{H\times W}\sum_{h=1}^{H}\sum_{w=1}^{W}\left[1-\text{Sim}[(m\cdot\Delta^{l}_{\epsilon_{\theta}})_{h,w},(m\cdot\Delta^{l}_{\epsilon_{\theta^{\prime}}})_{h,w}]\right],$ (S11)
@@
Textual Inversion [20]. We use the official implementation of Textual Inversion [^1] https://github.com/rinongal/textual_inversion, training it using stable diffusion v1.5. For each identity, the experiment is conducted with a batch size of 1 and a learning rate of 0.005 for 5,000 steps. The new token is initialized with the “person”.
Custom Diffusion [34]. We use the official implementation of Custom Diffusion [^2] https://github.com/adobe-research/custom-diffusion, and train it with stable diffusion v1.5. During training, the batch size is set to 1 and the learning rate to 1e-5. We generate 400 regularization images using stable diffusion version 1.5 with 50 steps of the DDIM sampler, prompted by the text “A photo of a person”. The model is trained for 300 training steps.
DreamBooth [66]. We use the third-party implementation of DreamBooth [^3] https://github.com/XavierXiao/Dreambooth-Stable-Diffusion, and train it with stable diffusion v1.5. Training is done by finetuning both the U-net diffusion model and the text transformer. The training batch size is 1 and the learning rate is set as 1e-6. The regularization images are generated with 50 steps of the DDIM sampler with the text prompt “A photo of a person”. For each identity, we train the model for 800 steps.
CelebBasis [88]. We adopt the official implementation of CelebBasis [^4] https://github.com/ygtxr1997/CelebBasis, and train it with stable diffusion v1.5. Training is conducted with a batch size of 2 for 800 steps. The learning rate is set to 0.005.
@@
Photomaker [40]. We use the official implementation and pre-trained models of Photomaker [^7] https://github.com/TencentARC/PhotoMaker, which is trained based on stable diffusion xl.
Our MasterWeaver. We employ stable diffusion v1.5 in our experiments, and our model is trained using a batch size of 16 and a learning rate of 1e-6. We initially train our method without $\mathcal{L}_{edit}$ and a face-augmented dataset for 100k iterations. Subsequently, we finetune it with $\mathcal{L}_{edit}$ and face augmented dataset for an additional 20k iterations $\lambda_{edit}$ is set as 0.01 and $\lambda_{disen}$ is set as 1. To enable classifier-free guidance, we use a probability of 0.05 to drop text and image individually, and a probability of 0.05 to drop text and image simultaneously. All experiments are conducted on 4$\times$×A800 GPUs with AdamW [47] optimizer.
@@
To evaluate the one-shot personalization, we randomly select 300 identities from the CelebA-HQ dataset [35], and each identity has one image as the reference. For quantitative analysis, we employ 50 prompts that encompass a range of clothing, styles, attributes, actions, and backgrounds. The full prompts are listed in Table S9. Following [40], we also collect a dataset consisting of 25 identities listed in Table S5, and each identity contains four images for evaluation under a multi-reference image setting. During evaluation, the identity embedding for FastComposer is computed as the mean embedding of the four images. For IP Adapter, the identity feature is constructed by concatenating the features from all four images, similar to our approach.
@@
Figure S9: Ablation on the value of $\lambda$.. As $\lambda$ increases, the face similarity, CLIP-I and DINO improve, yet slightly decreases the text alignment (CLIP-T).
@@
We have evaluated the impact of layer selection when calculating the editing direction loss $\mathcal{L}_{edit}$. From Table S6, incorporating Upblock2 and Upblock3 into the loss computation offers marginal improvements in editability at the cost of a significant decrease in identity consistency. Therefore, we only utilize Upblock1 for loss calculation in our implementation.
@@
To evaluate the effect of hyper-parameter $\lambda$ during inference, we vary its value from 0.1 to 1, and the results are illustrated in Fig. S9 and Fig. S10. It is observed that an increase in $\lambda$ enhances identity fidelity. Therefore, we keep $\lambda=1$ during generation for better identity consistency. However, users may choose to reduce $\lambda$ slightly to achieve better text alignment, at the cost of a minor reduction in identity fidelity.
@@
S9 More Quantitative Results
@@
Our method is primarily trained to generate images with a single face, and it may struggle with generating images consisting of multiple personalized identities. As illustrated in Fig. S18, directly combining two different identities fails to yield the intended results, instead producing a pair of faces that appear to be blends of the two. Utilizing techniques, such as attention rectification, can mitigate this issue, but a pre-defined location map for each identity involved is needed. In future work, we will explore more sophisticated approaches for the personalization of multiple identities. Furthermore, our method’s capability to generate photo-realistic images of specific individuals carries ethical considerations. Such a technique might be used for deepfake generation, particularly for offensive content. We will explore additional safeguarding methods to avoid this, such as integrating digital watermarks.

### v3

@@ 3.4 Learning Objective
Following [41], we employ a background disentanglement loss to encourage the identity feature only controlling the generation of facial region. Specifically, when the identity feature changes, we regularize the background to be unchanged:
@@
where $f=M\circ\psi(x)$ is the extracted identity feature and $1-m$ denotes the mask of background region. $A(\cdot)$ denotes the augmentation operation.
@@ 4.1 Experimental Settings
Evaluation dataset. To evaluate the one-shot personalization, we randomly select 300 identities from the CelebA-HQ dataset [37], and each identity has one image as the reference. Besides, our method can be extended for personalized generation using multiple reference images without additional training. Here, we also evaluate our method with multiple images as the reference. Following [42], we collect a dataset consisting of 25 identities, and each identity contains four images for evaluation. For quantitative analysis, we employ 50 prompts encompassing a range of clothing, styles, attributes, actions, and backgrounds. We randomly generate five images for each identity-prompt pair. More details are provided in Suppl. For visual comparison, we utilize the identity images collected from existing methods for a convenient comparison. Unless specifically mentioned, all images presented in this paper are produced using a single reference image.
Implementation details. We employ SD V1.5 in our experiments, and our model is trained using a batch size of 16 and a learning rate of 1e-6. To calculate the $\mathcal{L}_{edit}$, we utilize the output features of the cross attention blocks in the first decoder layer. $\lambda_{edit}$ is set as 0.01 and $\lambda_{disen}$ is set as 1. To enable classifier-free guidance, we use a probability of 0.05 to drop text and image individually, and a probability of 0.05 to drop text and image simultaneously. All experiments are conducted on 4$\times$×A800 GPUs with AdamW [50] optimizer. During image generation, we use 50 steps of the LMS sampler, and set the scale of classifier-free guidance as 5. More details of our model and implementation are provided in the Suppl.
@@
Figure 7: Ablation Study. The proposed losses and augmented dataset (denotes as aug) significantly improve the editability of the model. Besides, the identity fidelity consistently improves as the number increases.
@@ 4.2 Qualitative Comparison
Fig. 5 presents visual comparisons with existing methods using a single image as a reference. As depicted, MasterWeaver can generate photo-realistic personalized images in diverse scenarios, including modifications to attributes, clothing, background, and style. Optimization-based methods, i.e., Textual Inversion, Dreambooth, and Custom Diffusion, suffer from the overfitting in this limited-data scenario, leading to limited text alignment and compromised identity fidelity (e.g., rows 1$\sim$3). Tuning-free methods, i.e., Fastcomposer and IP Adapter maintain high identity fidelity but fall short in editing style and facial attributes (e.g., face attributes in rows 1 and 3, and the style in the last two rows). Photomaker shows improved text controllability; however, it sometimes fails to generate faithful identity (e.g., rows 1, 2, and 5). In comparison, our MasterWeaver can generate personalized images that are faithful to both reference identity and text prompts. Fig. 1 illustrates more results generated by our method. MasterWeaver can generate photo-realistic images with diverse clothing, accessories, facial attributes, and actions. Additionally, it allows for simultaneously editing multiple attributes.
MasterWeaver can also be directly extended for personalized image generation with multiple reference images of the same identity. Specifically, during inference, we simply concatenate the identity features of different images along the token dimension. As shown in Fig. 7, using more reference images improves the identity fidelity of the generated personalized images. Furthermore, we compare our method with competing methods in the setting with multiple reference images (i.e., with four reference images). From Fig. 6, one can see that MasterWeaver still outperforms competitors in identity fidelity and text controllability. More qualitative results can be found in Suppl.
Table 3: Ablation Study. The proposed background disentanglement loss $L_{disen}$, editing direction loss $L_{edit}$, and the face-augmented dataset (denote as aug) can improve the editability of MasterWeaver, while keeping the identity fidelity.
@@ 4.3 Quantitative Comparison
In addition to the qualitative comparisons, we further conduct the quantitative evaluation to validate the performance of MasterWeaver. Table 1 reports the comparisons conducted with a single reference image. From the table, we see that MasterWeaver outperforms existing state-of-the-art methods regarding text alignment and face similarity, demonstrating its ability to generate personalized images that maintain consistent identity and good alignment with texts. Moreover, MasterWeaver exhibits a competitive inference speed, requiring only 4s to generate an image on a single V100 GPU. Table 2 reports similar findings in a scenario utilizing four reference images, further demonstrating the effectiveness of MasterWeaver. While MasterWeaver’s scores for CLIP-I and DINO are not the highest, it should be noted that these metrics are not specifically designed for facial analysis and can be affected by extraneous factors such as pose. Additionally, we have conducted a user study to compare MasterWeaver with other methods. The detailed results are provided in Suppl.
@@ 4.4 Ablation Study
We have conducted ablation studies to evaluate the roles of various components in our method, including the background disentanglement loss $\mathcal{L}_{disen}$, editing direction loss $\mathcal{L}_{edit}$, and the face-augmented dataset.
Effect of the disentanglement loss $\mathcal{L}_{disen}$. We first evaluate the impact of $\mathcal{L}_{disen}$. As illustrated in Fig. 7, without $\mathcal{L}_{disen}$, the model fails to generate images with desired clothing. This suggests that $\mathcal{L}_{disen}$ could ease the influence of identity information on the background, thereby enhancing the editability of the generated images. From Table 3, both text alignment and face similarity metrics decrease without $\mathcal{L}_{disen}$, demonstrating its effectiveness.
Effect of the editing direction loss $\mathcal{L}_{edit}$. We then assess the effect of $\mathcal{L}_{edit}$. As shown in Fig. 7, without $\mathcal{L}_{edit}$, MasteWeaver fails to generate the identities with proper facial attributes (i.e., old woman). From Table 3, we see that the CLIP-T score drops significantly without $\mathcal{L}_{edit}$. Though the preservation of identity drops slightly, it is acceptable.
Effect of the face-augmented dataset. We further evaluate the effect of our constructed face-augmented dataset. Fig. 7 demonstrates that the model trained with the face-augmented dataset exhibits improved editability (e.g., with the augmented dataset, MasterWeaver successfully generates the personalized image with a smiling expression). The results in Table 3 further confirm its effectiveness, as the text alignment score increases when using the augmented dataset. More ablations are provided in Suppl.
@@ 5 Conclusion
In this work, we proposed MasterWeaver, a novel tuning-free method capable of generating personalized images with high efficiency, faithful identity, and flexible editability. The proposed editing direction loss and face-augmented dataset significantly improved the model’s editability while maintaining identity fidelity. Extensive experiments have demonstrated that MasterWeaver outperformed the state-of-the-art methods and generated photo-realistic images that are faithful to both identity and texts. Our method was versatile for various applications, including personalized digital content creation and artistic endeavors. Moreover, the proposed editing direction loss had the potential to be applied to other domains (e.g., animals and objects), enlarging its applicability.
Acknowledgement. This work was supported in part by the National Key R&D Program of China (2021YFF0900500), and the National Natural Science Foundation of China (NSFC) under Grant No. 62441202, and the Hong Kong RGC RIF grant (R5001-18).
@@
Fig. S8 illustrates the detailed framework of MasterWeaver. As shown in the figure, the ID Mapper consists of a stack of cross attention, feed forward layer and self attention layer. To effectively incorporate the identity feature $f$ with text feature to steer the personalized generation, we adopt the dual cross attention mechanism [90, 85]. As illustrated in Fig. S8, for each cross attention block of SD, we further introduce two learnable projection layers $W^{f}_{K}$ and $W^{f}_{V}$. The identity feature is then integrated by $\text{Attention}(Q,K^{f},V^{f})$, where $K^{f}=W^{f}_{K}\cdot f$ and $V^{f}=W^{f}_{V}\cdot f$ represent the projected key and value matrices of the identity information, respectively. To fuse the identity information with the text information, we sum them by:
@@
where $\lambda$ is a trade-off hyperparameter and set as $1$ during training. During training, we optimize the parameters of ID Mapper and the projection layers simultaneously, and keep the parameters of SD fixed.
@@
To calculate the editing direction loss $\mathcal{L}_{edit}$, we utilize the output features from the cross attention blocks in the first decoder layer. Specifically, given the source prompt $y$, target prompt $y^{\prime}$ and reference image $x$, the editing directions of the $l$-th cross attention block for SD and MasterWeaver are $\Delta^{l}_{\epsilon_{\theta}}(y,y^{\prime})\in\mathbb{R}^{H\times W\times D}$ and $\Delta^{l}_{\epsilon_{\theta^{\prime}}}(y,y^{\prime},x)\in\mathbb{R}^{H\times W\times D}$, where $H$, $W$, and $D$ represent the height, width, and dimension of the output feature, respectively. Then, the editing direction loss is calculated as follows:
@@
$\mathcal{L}^{l}_{edit}=\frac{1}{H\times W}\sum_{h=1}^{H}\sum_{w=1}^{W}\left[1-\text{Sim}[(m\cdot\Delta^{l}_{\epsilon_{\theta}})_{h,w},(m\cdot\Delta^{l}_{\epsilon_{\theta^{\prime}}})_{h,w}]\right].$ (S11)
@@
Textual Inversion [20]. We use the official implementation of Textual Inversion [^1] https://github.com/rinongal/textual_inversion, training it using SD V1.5. For each identity, the experiment is conducted with a batch size of 1 and a learning rate of 0.005 for 5,000 steps. The new token is initialized with the word “person”.
Custom Diffusion [36]. We use the official implementation of Custom Diffusion [^2] https://github.com/adobe-research/custom-diffusion, and train it with SD V1.5. During training, the batch size is set to 1 and the learning rate to 1e-5. We generate 400 regularization images using SD V1.5 with 50 steps of DDIM sampling, prompted by the text “A photo of a person”. The model is trained for 300 steps.
DreamBooth [69]. We use the third-party implementation of DreamBooth [^3] https://github.com/XavierXiao/Dreambooth-Stable-Diffusion, and train it with SD V1.5. Training is done by finetuning both the U-net diffusion model and the text transformer. The training batch size is 1 and the learning rate is set as 1e-6. The regularization images are generated with 50 steps of the DDIM sampler with the text prompt “A photo of a person”. For each identity, we train the model for 800 steps.
CelebBasis [91]. We adopt the official implementation of CelebBasis [^4] https://github.com/ygtxr1997/CelebBasis, and train it with SD V1.5. Training is conducted with a batch size of 2 for 800 steps. The learning rate is set to 0.005.
@@
Photomaker [42]. We use the official implementation and pre-trained models of Photomaker [^7] https://github.com/TencentARC/PhotoMaker, which is trained based on SDXL.
Our MasterWeaver. We employ SD V1.5 in our experiments, and our model is trained using a batch size of 16 and a learning rate of 1e-6. We initially train our model without $\mathcal{L}_{edit}$ and the face-augmented dataset for 100k iterations. Subsequently, we finetune it with $\mathcal{L}_{edit}$ and the face augmented dataset for an additional 20k iterations. $\lambda_{edit}$ is set as 0.01 and $\lambda_{disen}$ is set as 1. To enable classifier-free guidance, we use a probability of 0.05 to drop text and image individually, and a probability of 0.05 to drop text and image simultaneously. All experiments are conducted on 4$\times$×A800 GPUs with AdamW [50] optimizer.
@@
To evaluate the one-shot personalization, we randomly select 300 identities from the CelebA-HQ dataset [37], and each identity has one image as the reference. For quantitative analysis, we employ 50 prompts that encompass a range of clothing, styles, attributes, actions, and backgrounds. The full prompts are listed in Table S10. Following [42], we collect a dataset consisting of 25 identities listed in Table S4, and each identity contains four images for evaluation under the multi-reference image setting. During evaluation, the identity embedding for FastComposer is computed as the mean embedding of the four images. For IP Adapter, the identity feature is constructed by concatenating the features from all the four images, similar to our approach.
@@
Table S5: Quantitative comparison on style- and attribute-irrelevant prompts under single reference setting. The best result is shown in bold, and the second best is underlined.
Custom Diffusion [36] 0.209 0.751 0.652 0.635 548
FastComposer [87] 0.210 0.755 0.624 0.651 3
IPAdapter [90] 0.212 0.752 0.649 0.604 3
PhotoMaker [42] 0.231 0.710 0.531 0.594 10
InstantID [82] 0.205 0.758 0.669 0.682 21
Ours 0.231 0.766 0.665 0.686 4
Ours w/o $\mathcal{L}_{edit}$ 0.216 0.768 0.662 0.688 4
@@
Figure S9: More ablation results on $\mathcal{L}_{edit}$.
Figure S10: Ablation on the value of $\lambda$. As $\lambda$ increases, the face similarity, CLIP-I and DINO improve, yet slightly decreases the text alignment (CLIP-T).
@@
Effects of $\mathcal{L}_{edit}$ on metrics. As shown in Fig. S9, $\mathcal{L}_{edit}$ could improve the editability (e.g., hair, expression, and style) while keeping the identity fidelity. However, the CLIP-I and DINO metrics are calculated by comparing generated faces with reference faces. As for prompts editing style and attribute, the scores for successfully generated images can be lower than those of simply ID-preserved images. Table S5 further reports the metrics calculated on style- and attribute-irrelevant prompts. One can see that our $\mathcal{L}_{edit}$ could improve text editability, while slightly affects the identity fidelity, demonstrating its effectiveness.
@@
We have evaluated the impact of layer selection when calculating the editing direction loss $\mathcal{L}_{edit}$. From Table S6, incorporating Upblock2 and Upblock3 into the loss computation brings marginal improvements in editability at the cost of a significant decrease in identity consistency. Therefore, we only utilize Upblock1 for loss calculation.
@@
To evaluate the effect of hyper-parameter $\lambda$ during inference, we vary its value from 0.1 to 1, and the results are illustrated in Fig. S10 and Fig. S11. It is observed that increasing $\lambda$ enhances identity fidelity. Therefore, we keep $\lambda=1$ during generation for better identity consistency. However, users may choose to reduce $\lambda$ slightly to achieve better text alignment, at the cost of a minor reduction in identity fidelity.
@@
S9 More Comparisons
S9.1 Comparison with InstantID
Fig. S13 illustrates the comparisons between our method and InstantID [82]. One can see that InstantID exhibits limited flexibility in facial poses and text editability. In contrast, our method demonstrates better text alignment. Fig. S13 left also shows the metrics of identity preservation and text editability. Our method shows better trade-off between identity and editability.
S9.2 More Qualitative Results
@@
Figure S13: Comparison with Instant ID.
S9.3 More Quantitative Results
[…]
Table S7: User study. The numbers indicate the percentage (%) of volunteers who favor the results of our method over those of the competing methods.
Text Fidelity 68.8 61.5 62.2 70.2 52.1
Identity Fidelity 64.0 70.1 55.3 53.5 66.8
Image Quality 58.2 74.1 68.8 71.8 54.5
User Study. We invite volunteers to perform a user study on the personalized T2I results od MasterWeaver and its competitors. Given an identity image, a text prompt, and two synthesized images (ours v.s. competitor’s), the participants are asked to select the better one from three aspects. i) Text Fidelity: Which image is more faithful to the input text prompt? ii) Identity Fidelity: Which image is more similar to the input person’s identity? iii) Image Quality: Which generated image shows better quality?
For each evaluation aspect, we employ 40 participants, and each participant is asked to evaluate 50 randomly selected sample images, i.e., 2000 responses in total. As shown in Table S7, our method achieves a better trade-off between text editability and identity fidelity. While our method receives similar preference to Photomaker regarding text fidelity, our identity fidelity is much better, demonstrating its superiority.
@@
Our method is primarily trained to generate images with a single identity, and it may struggle with generating images consisting of multiple personalized identities. As illustrated in Fig. S20, directly combining two different identities fails to yield the intended results, instead producing a pair of faces that appear to be blends of the two. Utilizing techniques, such as attention rectification, can mitigate this issue, but a pre-defined location map for each identity is needed. In future work, we will explore more sophisticated approaches for the personalization of multiple identities. Furthermore, our method’s capability to generate photo-realistic images of specific individuals carries ethical considerations. Such a technique might be used for deepfake generation, particularly for offensive content. We will explore additional safeguarding methods to avoid this, such as integrating digital watermarks.